A sodium hydroxide solution containing 40% by weight of pure NaOH has a specific gravity of 1.5. What volume of this solution will be required in the preparation of 500ml of a 0.045M NaOH solution 
 

Grade:11

1 Answers

Arun
4 years ago

Specific gravity=Density of substanceDensity of water
Since density of water is 1.We get specific gravity =Density of substance
Conder volume to b v.....Now Mass of naoh=1.5 v * (40/100)
Moles of naoh=Mass/40
Molaity=(moles/Volume)*10000
now


Mf Vf=MVi
((1.5*40)/40*100)*1000*Vfi=0.45*500
Vi=15ml